how do you find the points of intersection for
x^2 + y^2=25
x + y^2=5

Answer by edjones(8007)

x^2 + y^2=25
x + y^2=5
.
x=5-y^2
(5-y^2)^2+y^2=25 substitute 5-y^2 for x.
y^4-10y^2+25+y^2=25
y^4-9y^2=0
y^2(y^2-9)=0
y^2(y+3)(y-3)=0
y=0, x=5; y=3, x=-4; y=-3 x=4
.
Ed
.